OpenStudy (anonymous):

find the slope and y-int. of the given points: (-8, .6), (2,-2.4). show all steps

OpenStudy (aroub):

-2.4-.6/ 2+8 -3/10 y=mx+b , m=-3/10 , (2,-2.4) -2.4=-3/10(2)+b -2.4=-6/10+b -2.4=-3/5+b 3/5-12/5=b -9/5=b y=-3/10x-9/5

OpenStudy (aroub):

slope=-3/10

OpenStudy (aroub):

y-intercept= -9/5

OpenStudy (anonymous):

thank you!